# OThink-MR1: Stimulating multimodal generalized reasoning capabilities via dynamic reinforcement learning

### v1

@@ 4.2.1. Cross Task Overall Performance
We validate the multimodal generalization ability on multimodal content understanding and reasoning tasks, i.e., visual counting(VC) and geometry reasoning(GR) tasks respectively. The multimodal reasoning tasks are more challenging than multimodal content understanding as they require complex logical deduction beyond direct multimodal interpretation. To assess generalization, we conduct cross-task experiments in both directions:
@@
From the results shown in Figure 7, we can observe that:
@@
  • In the Reasoning to Understanding Generalization experiment, the model with SFT post-training shows similar performance to the model without post-training In the Understanding to Reasoning Generalization experiment, the model with SFT experiences a -77.77% performance drop. These results indicate that SFT tends to memorize training data rather than learn generalized knowledge, leading to poor cross-task generalization, especially in the more challenging Understanding to Reasoning scenario.
  • In contrast, the model with GRPO-D post-training achieves significant improvements: a +58.82% relative gain in Reasoning to Understanding and a +3.09% relative gain in Understanding to Reasoning compared to the model without post-training. These results demonstrate that GRPO-D effectively enables the model to acquire task-generalized knowledge in complex multimodal tasks.
These results are promising, as they demonstrate that a model post-trained with GRPO-D on one task can effectively generalize to other multimodal tasks. Whether post-trained on simple content understanding or complex reasoning tasks, GRPO-D consistently achieves strong generalization, especially outperforming SFT. highlights GRPO-D’s ability to foster genuine understanding and reasoning of multimodal knowledge, while SFT is limited to memorization.
@@ 4.2.2. Case Study
To further evaluate the model’s generalized reasoning ability, we perform case studies on cross-task generalization. We select representative samples from both Reasoning to Understanding and Understanding to Reasoning experiments, as shown in Figure 8.
From Figure 8(a), we observe that after GRPO-D post-training on the geometry reasoning task, the model generalizes effectively to detect and recognize objects in the image, such as identifying the green bus and blue motorcycle. Similarly, in Figure 8(b), GRPO-D post-training on the visual counting task enables the model to analyze relationships within visual data. A notable example is the model’s output, which incorporates the inscribed angle theorem–a key factor contributing to the correct result. These results demonstrate GRPO-D’s strong generalized reasoning ability, as evidenced by its adaptable and logical thinking process across diverse tasks.
4.3. Same-Task Evaluation (RQ3)
In Othiner-MR1, we introduce GRPO-D, which incorporates a dynamic KL divergence strategy to better balance exploration and exploitation. Then, we present the following empirical results to demonstrate its effectiveness in same-task evaluation. The overall performance on the two adopted datasets for same-task validation is presented in Table 1. For Qwen2-VL-2B-Instruct+GRPO, we use the optimal hyperparameters: $\alpha=1$, a constant $\beta(s)=0.01$ for geometry reasoning task and $\alpha=1$, a constant $\beta(s)=0.04$ for visual counting task.
Model Accuracy
VC GR
Qwen2-VL-2B-Instruct 42.50% 15.52%
Qwen2-VL-2B-Instruct+GRPO 64.50% 30.24%
Qwen2-VL-2B-Instruct+SFT 68.50% 32.49%
Qwen2-VL-2B-Instruct+GRPO-D 76.50% 34.35%
From the results, we can observe that:
  • Qwen2-VL-2B-Instruct+GRPO underperforms Qwen2-VL-2B-Instruct+SFT. This is primarily because, with sufficient training data, SFT can effectively memorize the comprehensive data distribution, leading to better performance in same-task validation.
  • Qwen2-VL-2B-Instruct+GRPO-D outperforms Qwen2-VL-2B-Instruct+GRPO with constant KL divergence: The dynamic KL strategy in GRPO-D achieves relative improvements of +18.60% and +13.59% for visual counting and geometry reasoning tasks, respectively. This improvement is attributed to the dynamic KL strategy’s ability to balance exploration and exploitation more effectively.
  • Qwen2-VL-2B-Instruct+GRPO-D outperforms Qwen2-VL-2B-Instruct+SFT: Compared to SFT post-training, GRPO-D achieves relative improvements of +11.68% and +5.72% on the same tasks, further demonstrating the effectiveness of the dynamic KL strategy in enhancing model performance.
@@ 5. Conclusion and Discussion
In this work, we extend the GRPO to MLLM and make several insightful discoveries, as detailed below:
  • For same-task validation, GRPO greatly outperforms than SFT in early training stages. However, as training steps increase, SFT matches or surpasses GRPO. This phenomenon occurs because SFT, when exposed to sufficient training data, can memorize the comprehensive distribution of the data. GRPO, on the other hand, struggles with balancing exploration and exploitation, leading to suboptimal performance. To address this, we propose a dynamic KL divergence mechanism, enabling GRPO to consistently outperform SFT.
  • When trained and validated on different task types (e.g., multimodal understanding vs. reasoning), GRPO-D significantly outperforms SFT and the base model without post-training. This demonstrates GRPO-D’s great generalized ability to enable MLLMs to achieve deeper understanding and reasoning, highlighting its potential for developing generalized MLLMs applicable to diverse multimodal tasks.
In conclusion, our work proposes a novel framework, OThink-MR1, which incorporates dynamical reinforcement learning for multimodal language language models. This framework enables the development of generalizable multimodal models, representing a significant advancement in the field of multimodal generalized reasoning.

### v2

@@ 4.3.1. Cross Task Overall Performance
We validate the multimodal generalization ability on multimodal content understanding and reasoning tasks, i.e., visual counting(VC) and geometry reasoning(GR) tasks respectively. The multimodal reasoning tasks are more challenging than multimodal content understanding as they require complex logical deduction beyond direct multimodal interpretation. To assess generalization, we conduct cross-task experiments in both directions as follows:
@@
The overall cross-task validation results are presented in Figure 7 and we can observe that:
@@
  • In the Reasoning to Understanding Generalization experiment, the model with SFT post-training shows similar performance to the model without post-training In the Understanding to Reasoning Generalization experiment, the model with SFT experiences a relative -77.77% performance drop. These results indicate that SFT tends to memorize training data rather than learn generalized knowledge, leading to poor cross-task generalization, especially in the more challenging Understanding to Reasoning scenario.
  • In contrast, the model with GRPO-D post-training achieves significant improvements in both generalization experiments. For the Reasoning to Understanding Generalization experiment, it gains a +63.53% (vs. without post-training) and +61.63% (vs. SFT) relative improvement. For Understanding to Reasoning Generalization experiment, it gains a +4.25% (vs. without post-training) and +368.99% (vs. SFT) relative improvement. These results demonstrate that GRPO-D effectively enables the model to acquire task-generalized knowledge in complex multimodal tasks, particularly compared with SFT post-training method.
These results are promising, as they demonstrate that a model post-trained with GRPO-D on one task can effectively generalize to other multimodal tasks. This is particularly evident when the model is post-trained on complex reasoning task. Overall, these results highlight GRPO-D’s ability to foster genuine understanding and reasoning of multimodal knowledge, while SFT is often limited to memorization.
@@ 4.3.2. Case Study
To further evaluate the model’s generalized reasoning ability, we perform case studies on cross-task generalization. We select representative samples from both Reasoning to Understanding and Understanding to Reasoning experiments, as shown in Figure 8. We can observe that after GRPO-D post-training on the geometry reasoning task, the model generalizes effectively to detect and recognize objects in the image, such as identifying the green bus and blue motorcycle. Similarly, in Figure 8(b), GRPO-D post-training on the visual counting task enables the model to analyze relationships within visual data, such as incorporating the inscribed angle theorem—a key factor contributing to the correct result. These results demonstrate GRPO-D’s strong generalized reasoning ability, as evidenced by its adaptable and logical thinking process across diverse tasks.
@@ 5. Conclusion and Discussion
In this work, we propose OThink-MR1, an advanced MLLM equipped with profound comprehension and reasoning capabilities across multimodal tasks. The key contributions are detailed below:
  • For same-task validation, GRPO initially outperforms SFT in early training stages. However, as training steps increase, SFT matches or surpasses GRPO. This phenomenon occurs because SFT, when exposed to sufficient training data, can memorize the comprehensive distribution of the data. GRPO, on the other hand, struggles with balancing exploration and exploitation, leading to suboptimal performance. To address this, we propose a dynamic reinforcement learning algorithm, named GRPO-D with dynamic KL divergence. Extensive experiments demonstrate that GRPO-D consistently outperforms SFT.
  • When trained and validated on different task types (e.g., multimodal understanding vs. reasoning), GRPO-D significantly outperforms SFT and the base model without post-training. This demonstrates GRPO-D’s strong generalized ability to enable MLLMs to achieve deeper understanding and reasoning, highlighting its potential for developing generalized MLLMs applicable to diverse multimodal tasks.
Overall, OThink-MR1 enables the development of generalizable multimodal models, representing an advancement in the field of multimodal generalized reasoning. Additionally, several interesting findings have emerged that warrant further exploration: the optimal range of KL weight values appears to be correlated with the complexity of the tasks. Moreover, the generalization ability of GRPO-D appears to be influenced by the reasoning demands placed on the training samples. In the future, we plan to explore these relationships more deeply.